6. The decay constant of uranium -238 is 4.9 x 10-18s-1. Calculate its half life in seconds and in years.


1
Expert's answer
2021-03-24T19:42:30-0400

"T_{1/2}=\frac{\ln2}{\lambda}=\frac{\ln2}{4.9\cdot 10^{-18}}\approx141\cdot10^{15}\ (s)=4.486\cdot10^9 (years)" . Answer